OpenStudy (anonymous):

the preimeter of a rectangle is 38 if the length is 3 more than the width, find the width

OpenStudy (mathteacher1729):

Have you tried sketching a picture?

OpenStudy (anonymous):

Perimeter = 2L + 2W But, L = W +3 so now our equation is: P = 2(W+3) + 2W P = 2W +6 + 2W P = 4W + 6 We are told that P = 38 so... 38 = 4W +6 32 = 4W W = 8